OpenStudy (anonymous):

help step by step. 40n^3-20n^2+32n-16

OpenStudy (anonymous):

@jojo4eva you know how to do this?

OpenStudy (jojo4eva):

So divide each term by 4

OpenStudy (anonymous):

10n^3-5n^2 + 8n-4?

OpenStudy (anonymous):

sub in n as 1/2, u get 5-5+16-16=0, then u can troll and divide the eqn by (x-1/2), solve the quadratic with quadratic eqn. Bam Trolled

OpenStudy (anonymous):

or u can divide using 2x-1 if u are not comfortable with fraction division

OpenStudy (anonymous):

your still wrong bro your suppose to factor by grouping

OpenStudy (anonymous):

nvm, 帮人帮到低,送佛送到西, u get (2x-1)(5x^2-4)

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

so. 40n^3-20n^2+32n-16=4(2n-1)(5n^2-4), 5

OpenStudy (anonymous):

5n^2-4, thus the other two answer are (plus/minus) \[\sqrt{5}\div4\]